OpenStudy (anonymous):

A line passes through (9, –9) and (10, –5).Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

equation for the line passes through the point (x1,y1) and (x2,y2) is y-y1 = m (x-x1) where m = slope = (y2-y1)/ (x2-x1)